what is 44 minus twice a number is 11 more than the number
Answer by ewatrrr(24785)

Hi,
Let x represent the number
.
Question states: Write as You Read
44 - 2n = n +11
.
or
n + 11 = 44 -2n
3n + 11 = 44
3n = 33
n = 11